# SAKE: Towards Editing Auditory Attribute Knowledge of Large Audio-Language Models

### v1

@@ Abstract
Knowledge editing offers an efficient way to update model knowledge without full retraining, but prior work has concentrated almost exclusively on textual or visual modalities. We introduce SAKE, the first benchmark specifically designed for editing auditory attribute knowledge in Large Audio-Language Models (LALMs). Unlike factual updates, SAKE targets several abstract auditory attributes, capturing knowledge types that go beyond conventional textual and visual domains. We benchmark seven editing methods on two LALMs along four dimensions: reliability, generality, audio/text locality, and portability. Results highlight challenges such as preserving intra-attribute knowledge unrelated to the edit, generalizing edits to multimodal reasoning, and maintaining edits under sequential updates. SAKE provides a principled framework to study how knowledge editing extends to the auditory modalities, opening new directions for maintaining and adapting LALMs in more diverse real-world scenarios.
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023; Grattafiori et al., 2024; Hurst et al., 2024) have demonstrated remarkable progress across natural language understanding and generation tasks. As they scale, knowledge editing (De Cao et al., 2021; Mitchell et al., 2022; Zheng et al., 2023; Deng et al., 2025) has emerged as an important line of research. The goal of knowledge editing is to efficiently update specific pieces of model knowledge without full retraining while minimizing the risk of catastrophic forgetting. These techniques enable incorporating new facts (De Cao et al., 2021), correcting errors, mitigating biases (Chen et al., 2024a), and supporting personalization (Lu et al., 2025d), making knowledge editing a key tool for adapting LLMs to real-world use.
@@
Advances in LLMs have also led to multimodal extensions such as large vision-language models (LVLMs) (Liu et al., 2023; Li et al., 2023) and large audio-language models (LALMs) (Lu et al., 2025b; Chu et al., 2024; Kuan et al., 2024; Ghosh et al., 2025; Hurst et al., 2024; Lin et al., 2025; Lu et al., 2025a; Gong et al., 2023; Tang et al., 2024), which integrate additional modalities into LLMs. As multimodal models gain adoption, it becomes increasingly important to extend knowledge editing beyond the textual domain. Recent efforts have explored editing LVLMs through benchmarks targeting visual knowledge updates (Cheng et al., 2023; Huang et al., 2024; Zhang et al., 2024). However, knowledge editing in auditory modalities remains unexplored despite the rising importance of LALMs for speech and audio understanding.
@@
Figure 1: Overview of SAKE benchmark. Four auditory attributes are targeted. Reliability checks if the edit succeeds, Generality if it holds for equivalent data, Locality if unrelated knowledge remains unchanged, and Portability if it transfers to related knowledge. For example, after editing “frog” to “dog,” the answer of the portability question should change from “Insectivore” to “Omnivore.”
Editing auditory attribute knowledge (Yang et al., 2025a) introduces unique challenges and opportunities. Factual knowledge (Thorne et al., 2018) typically involves discrete and concrete statements (e.g., “Paris is the capital of France”). In contrast, auditory attributes such as speaker gender, emotion, spoken language, or animal sounds represent high-level and continuous perceptual concepts. These attributes may manifest through infinitely many acoustic realizations, including different speakers, prosodic variations, or recording conditions, while still grounding the same underlying attribute. As a result, it is unclear whether existing methods developed for editing discrete factual knowledge can be extended to such abstract representations. At the same time, editing auditory attributes can extend applications in the textual domain, such as debiasing (Chen et al., 2024a) and personalization (Lu et al., 2025d), into the auditory modalities, with new possibilities such as mitigating gender biases in LALMs (Lin et al., 2024b) and enabling personalization (Yang et al., 2025c; Lee et al., 2024) by adapting to a user’s unique voice or speaking style.
To this end, we introduce SAKE (Speech and Audio Attribute Knowledge Editing Benchmark), the first benchmark for auditory attribute knowledge editing in LALMs (Figure 1). SAKE covers four auditory attributes: speaker gender, speaker emotion, spoken language, and animal sounds, and evaluates editing methods along four dimensions (Yao et al., 2023): reliability (whether edits succeed), generality (whether they transfer to equivalent variations), locality (whether unrelated knowledge is preserved), and portability (whether edits propagate to interconnected knowledge).
Through comprehensive experiments on two strong LALMs, DeSTA2.5-Audio (Lu et al., 2025b) and Qwen2-Audio (Chu et al., 2024), we benchmark diverse editing techniques and reveal significant challenges in editing auditory attribute knowledge, such as preserving intra-attribute knowledge that is irrelevant to the edits and ensuring that updated auditory attributes propagate to reasoning over related knowledge. Moreover, when multiple edits are applied sequentially, most methods suffer from severe forgetting, often losing previously edited knowledge after a few additional updates. Overall, our contributions are two-fold [^1] Resources will be available at https://github.com/ckyang1124/SAKE:
  1. 1.
    To the best of our knowledge, this is the first work to explore knowledge editing in auditory modalities of LALMs, an area that has remained unexplored. We introduce the first comprehensive benchmark for auditory knowledge editing and evaluate seven common editing methods, providing novel insights into this direction.
  2. 2.
    Although these editing methods are effective in textual and visual domains, editing auditory attributes, which are abstract perceptual concepts, remains challenging. We show that current methods struggle both to generalize edited knowledge and to preserve unrelated knowledge, underscoring the need for future advances in auditory knowledge editing.
@@ 2.1 Knowledge Editing
Knowledge editing (Zhu et al., 2024) refers to techniques designed to efficiently update or modify the knowledge stored in models. The goal is to adjust model knowledge in a lightweight manner while avoiding catastrophic forgetting that may arise from retraining directly on the target knowledge. Existing methods (Mitchell et al., 2022; De Cao et al., 2021; Zheng et al., 2023; Meng et al., 2022) adopt various strategies: using a hypernetwork to predict parameter updates for incorporating new knowledge (De Cao et al., 2021; Mitchell et al., 2022), identifying and adjusting neurons associated with specific knowledge (Meng et al., 2022), or leveraging in-context learning (ICL) to enable models to acquire updated knowledge (Zheng et al., 2023). More recent works have extended the line of work from Meng et al. (2022) to mass editing (Meng et al., 2023) and unstructured scenarios (Deng et al., 2025; Jiang et al., 2025; Su et al., 2025). Beyond correcting factual knowledge, these techniques have also been applied to tasks such as bias mitigation (Chen et al., 2024a), detoxification (Wang et al., 2024a), personalization (Lu et al., 2025d), and unlearning (Li et al., 2025).
More recently, researchers have begun to investigate knowledge editing in large vision-language models (LVLMs)(Liu et al., 2023; Li et al., 2023). For example, Cheng et al. (2023) introduced the MMEdit benchmark to explore editing visual knowledge, while subsequent works such as VLKEB (Huang et al., 2024) and MC-MKE (Zhang et al., 2024) expanded the evaluation scope to provide a more comprehensive understanding of editing in visual modalities. However, no prior work has examined editing auditory attribute knowledge in LALMs, which involves abstract perceptual concepts rather than concrete facts, distinguishing our study from existing research.
@@ 2.2 Large Audio-Language Models (LALMs)
LALMs extend text-based LLMs to auditory modalities such as speech and audio, opening new possibilities for auditory understanding (Gong et al., 2023; Tang et al., 2024; Chu et al., 2023; 2024; Lu et al., 2025b; Ghosh et al., 2025). These models typically integrate auditory encoders (Radford et al., 2023) with an LLM backbone (Touvron et al., 2023; Yang et al., 2024a) through fine-tuning. While these works advance the integration of auditory knowledge into text-based LLMs, little attention has been given to how auditory-specific knowledge can be edited or updated, which motivates our study.
@@ 3.1 Problem Formulation
In this work, we focus on editing auditory attribute knowledge within LALMs, including their perception and understanding of speaker gender, emotion, spoken language, and animal sounds. In this setting, $y_{e}$ corresponds to new auditory attribute labels (e.g., emotions or languages) that differ from the original attribute labels $y_{o}$ associated with $a_{e}$. For example, given a speech labeled with a happy emotion, we may edit the LALM so that it instead perceives the speech as having a sad tone.
@@ 3.2 Evaluation Dimensions and Corresponding Metrics
The edited models should not only generate the correct edit target for the editing data itself but also produce consistent outputs for equivalent neighborhoods of the editing data, such as speech samples sharing the same emotion as $a_{e}$ or paraphrased variants of $x_{e}$. This requirement is quantified by the generality metric $S_{gen}$, defined as
@@
$S_{loc}=\mathbb{E}_{\begin{subarray}{c}(a_{e},x_{e},y_{e})\sim\mathcal{D}_{edit}\\ (a_{\ell},x_{\ell},y_{\ell})\sim\mathcal{L}(a_{e},x_{e},y_{e})\end{subarray}}\Bigl[\mathbb{I}\bigl(f(a_{\ell},x_{\ell};\theta^{\prime})=f(a_{\ell},x_{\ell};\theta)\bigr)\Bigr].$ (3)
@@
Knowledge is not completely disentangled or isolated but rather interconnected. Editing one piece of knowledge may influence other related knowledge. For example, if we edit an LALM’s perception of a frog’s sound to that of a dog, the model’s knowledge of the corresponding physical characteristics of that animal should also be updated. The portability metric $S_{port}$ evaluates how well the edited model generalizes the updated knowledge to other related knowledge:
@@
Here, $\mathcal{P}(a_{e},x_{e},y_{e})$ denotes the set of data connected to the edited knowledge, with $a_{p}$, $x_{p}$, and $y_{p}$ representing the auditory input, text input, and ground-truth labels of these connected instances.
@@ 3.3 Dataset Construction
Involved Auditory Attribute Knowledge and Audio Sources. SAKE focuses on editing the knowledge of four different auditory attributes: speaker gender, speaker emotion, spoken language, and animal sound. They are chosen for their importance in many downstream applications.
We source audio data from the SAKURA benchmark (Yang et al., 2025b), which evaluates LALMs on recognizing these four auditory attributes and related multi-hop reasoning. SAKURA compiles audio samples and attribute labels from CommonVoice (Ardila et al., 2020), CREMA-D (Cao et al., 2014), ESC-50 (Piczak, 2015), and the Animal-Sound Dataset (Şaşmaz & Tek, 2018). We also extract the auditory attribute labels from SAKURA for these attributes to form a set of labels [^2] Gender: Male, Female; Language: English, German, Spanish, French, Italian, Chinese, Japanese, Korean; Emotion: Happy, Disgust, Sad, Fear, Angry; Animal: Dog, Cat, Pig, Cow, Frog, Hen, Rooster, Sheep, Crow..
Editing Pair Creation. We begin by constructing editing pairs $(y_{o},y_{e})$, where $y_{o}$ is the original attribute label and $y_{e}$ the target label after editing. For each attribute, we generate 300 editing pairs by uniformly sampling one label as $y_{o}$ and another distinct label as $y_{e}$. For example, the editing pair $(\text{dog},\text{cat})$ represents an edit in which the model’s perception and understanding of dog sounds are updated to those of cat sounds. To avoid bias, we ensure that all labels for a given attribute appear with approximately equal frequency as both $y_{o}$ and $y_{e}$.
Reliability Dataset Construction. For each attribute and a given editing pair $(y_{o},y_{e})$, we sample an audio instance $a_{e}$ with label $y_{o}$ from the dataset, along with the corresponding text questions $x_{e}$ from the SAKURA benchmark that prompt the model to recognize the attribute. Together with the edit target $y_{e}$, this forms an editing instance $(a_{e},x_{e},y_{e})$. For each attribute, we construct 300 such instances, resulting in a total of 1,200 editing instances. This dataset, denoted as $\mathcal{D}_{edit}$, is used both for applying the edits and for evaluating the reliability of the editing methods.
@@
Generality Dataset Construction. For each editing instance $(a_{e},x_{e},y_{e})$ in $\mathcal{D}_{edit}$, we construct its equivalent neighborhood $\mathcal{N}(a_{e},x_{e})={(a_{e}^{\prime},x_{e}^{\prime})}$ by sampling an alternative audio $a_{e}^{\prime}$ with the same attribute label as $a_{e}$ from the dataset and by paraphrasing the text question $x_{e}$ into $x_{e}^{\prime}$. Based on these variations, we create testing instances for evaluating generality, considering three types of cases: Type 1: Equivalent neighborhood of the text modality $(a_{e},x_{e}^{\prime})$; Type 2: Equivalent neighborhood of the auditory modality $(a_{e}^{\prime},x_{e})$; Type 3: Equivalent neighborhood involving both auditory and text modalities, resulting in instances of the form $(a_{e}^{\prime},x_{e}^{\prime})$. By incorporating these three types of generality testing data, we comprehensively assess how well the editing methods extend the edited knowledge across the equivalent neighborhood.
@@
Locality Dataset Construction. Similar to the construction of the generality dataset, for each editing instance $(a_{e},x_{e},y_{e})$ in $\mathcal{D}_{edit}$, we construct an out-of-scope dataset $\mathcal{L}(a_{e},x_{e},y_{e})=\{(a_{\ell},x_{\ell},y_{\ell})\}$. When $a_{\ell}\neq\text{None}$, we consider four types of knowledge locality associated with auditory modalities, which we refer to as Audio Locality in this paper, as illustrated in Figure 1. First, editing one attribute should not affect others. We sample a question-answer pair from the SAKURA benchmark that requires the LALM to recognize attributes different from those in the editing instance as $(a_{\ell},x_{\ell},y_{\ell})$. Second, even within the same attribute, labels that are not involved in the edit (i.e., neither $y_{o}$ nor $y_{e}$) should not be perturbed. Accordingly, we sample a question-answer pair from SAKURA requiring recognition of the same attribute as the editing instance, but with ground truth differing from both $y_{o}$ and $y_{e}$. Note that for the gender attribute, which only has two labels, it is infeasible to construct such cases; therefore, this type of locality data is not included for gender. Third, when editing from $y_{o}$ to $y_{e}$, the model’s original knowledge of $y_{e}$ should be preserved, which we test with a question-answer pair from SAKURA that requiring recognizing the same attribute as the editing instance, with ground truth $y_{e}$. Finally, editing should not interfere with general auditory processing. We source question-answer pairs from Dynamic-SUPERB Phase-2 (Huang et al., 2025), a benchmark covering diverse auditory processing tasks. We exclude tasks involving the four attributes considered in our work to ensure irrelevance to the edited knowledge.
Conversely, to assess the preservation of purely text-based knowledge when $a_{e}=\text{None}$ (text locality), we use question–answer data from MMLU (Hendrycks et al., 2021).
@@
Portability Dataset Construction. For each editing instance $(a_{e},x_{e},y_{e})$ in $\mathcal{D}_{edit}$, we construct a set of connected knowledge $\mathcal{P}(a_{e},x_{e},y_{e})$ associated with the edited attribute. Since SAKURA is designed to evaluate how well LALMs integrate internal world knowledge with auditory attribute knowledge, it already provides relevant knowledge linked to auditory attributes (e.g., physical characteristics of animal labels). Building on this, we create questions that specifically target $\mathcal{P}(a_{e},x_{e},y_{e})$. To avoid ambiguity, we ensure that the answers to these questions are not simultaneously valid for both $y_{o}$ and $y_{e}$. For example, if $y_{o}$ is dog and $y_{e}$ is cat, a question about the animal’s physical characteristics will not use an answer like “tail,” which applies to both. This guarantees that the knowledge examined in the portability dimension indeed requires updating after the edit.
@@
Training Dataset Construction and the Dataset Summary. We also prepare a separate training dataset to accommodate editing methods that require (1) training on auxiliary data or (2) access to additional data beyond the testing instances. Its construction follows the same procedures as those used for the reliability, generality, and locality datasets, with training and testing sets kept fully disjoint to avoid leakage. In total, the training set contains 4,000 instances (32,000 speech/audio files and 36,000 QA pairs), while the testing set comprises 1,200 instances (10,800 speech/audio files and 12,000 QA pairs). A summary of dataset statistics is provided in Appendix B.

### v2

@@ Abstract
Knowledge editing enables targeted updates without retraining, but prior work focuses on textual or visual facts, leaving abstract auditory perceptual knowledge underexplored. We introduce SAKE, the first benchmark for editing perceptual auditory attribute knowledge in large audio-language models (LALMs), which requires modifying acoustic generalization rather than isolated facts. We evaluate eight diverse editing methods on three LALMs across reliability, generality, locality, and portability, under single and sequential edits. Results show that most methods enforce edits reliably but struggle with auditory generalization, intra-attribute locality, and multimodal knowledge propagation, and often exhibit forgetting or degeneration in sequential editing. Additionally, fine-tuning the modality connector emerges as a more robust and balanced baseline compared with directly editing the LLM backbones. SAKE reveals key limitations of current methods and provides a foundation for developing auditory-specific LALM editing techniques.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023; Grattafiori et al., 2024; Hurst et al., 2024) have achieved remarkable success across a wide range of language tasks. As they scale and are increasingly deployed, knowledge editing (De Cao et al., 2021; Mitchell et al., 2022; Zheng et al., 2023; Deng et al., 2025) has become a key technique for efficiently updating specific model knowledge without full retraining while minimizing catastrophic forgetting. Knowledge editing can also be extended to bias mitigation (Chen et al., 2024a) and personalization (Lu et al., 2025d).
@@
Recent advances have extended LLMs to multimodal settings, including large vision-language models (LVLMs) (Liu et al., 2023; Li et al., 2023) and large audio-language models (LALMs) (Lu et al., 2025b; Chu et al., 2024; Kuan et al., 2024; Ghosh et al., 2025b; Hurst et al., 2024; Lin et al., 2025; Lu et al., 2025a; Tang et al., 2024). As multimodal models become more prevalent, it is increasingly important to understand whether and how knowledge editing can be applied beyond the textual domain. While recent benchmarks have explored visual knowledge editing in LVLMs (Cheng et al., 2023; Huang et al., 2024; Zhang et al., 2025), knowledge editing in the auditory modality remains largely unexplored.
Most knowledge editing studies (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022) focus on factual knowledge expressed as declarative propositions, often formalized as subject-relation-object triplets (Meng et al., 2022) (e.g., “Paris is the capital of France”). Visual knowledge editing in LVLMs largely follows this paradigm by treating images as alternative grounding for entity-centric facts. In contrast, perceptual knowledge is qualitatively different. Editing perceptual knowledge is conceptually closer to modifying style-level understanding, such as writing style in the text domain or image style in the vision domain, than to updating individual facts, as it operates on high-level abstractions rather than isolated factual associations.
Auditory attribute knowledge is a form of perceptual knowledge that reflects how models conceptualize attributes such as speaker gender and emotion. These attributes arise from perceptual abstractions over signals with substantial intra-class variability across speakers, prosody, and recording conditions. Editing such knowledge, therefore, requires modifying how models generalize perceptual evidence, rather than changing a factual statement. Hence, it is unclear whether editing methods developed for factual knowledge can generalize to abstract perceptual concepts in auditory modalities.
@@
Figure 1: Left: Conceptual illustration of the knowledge scope affected by an edit. Right: Overview of SAKE benchmark, targeting auditory attributes including speaker gender, emotion, spoken language, and animal sounds. Reliability measures edit success, Generality its consistency across equivalent data, Locality the preservation of unrelated knowledge, and Portability its transfer to connected knowledge. For example, after editing “frog” to “dog,” the answer of the portability question should change from “Insectivore” to “Omnivore.”
To this end, we introduce SAKE (Speech and Audio Attribute Knowledge Editing Benchmark), the first benchmark for auditory attribute knowledge editing in LALMs (Figure 1). Auditory attribute knowledge editing targets modifications of an LALM’s perceptual understanding of specific attributes, such as changing its auditory knowledge of frogs’ vocalizations to that of dogs (Figure 1), while preserving related reasoning and unrelated knowledge. Such editing enables applications such as LALM error correction for misrecognition, debiasing, and voice assistant personalization, which are inherently more challenging than text-based settings due to the complexity of auditory modalities. SAKE spans diverse auditory attributes and evaluates editing methods along four dimensions (Yao et al., 2023): reliability, generality, locality, and portability.
We evaluate eight representative editing methods on three LALMs, DeSTA2.5-Audio (Lu et al., 2025b), Qwen2-Audio (Chu et al., 2024), and Audio Flamingo 3 (Ghosh et al., 2025a). While these methods have proven effective in textual domains, our results reveal persistent challenges in auditory attribute knowledge editing, including preserving intra-attribute knowledge and enabling edit propagation to connected reasoning. Under sequential editing, most methods suffer severe forgetting. Together, these findings highlight the unique challenges of auditory knowledge editing and the need for future advances in this research area.
Our contributions are three-fold: (1) We introduce SAKE, the first comprehensive benchmark for systematically evaluating auditory attribute knowledge editing of LALMs; (2) We show that existing methods struggle to preserve non-target auditory knowledge and to propagate edits during reasoning, revealing key limitations in editing abstract perceptual attributes; (3) Through extensive analysis, we show that fine-tuning the modality connector is strong and more balanced across the evaluation dimensions, highlighting an important direction for auditory-specific editing techniques.
@@ 2.1 Knowledge Editing
Knowledge editing (Zhu et al., 2024) aims to efficiently modify the models’ knowledge while avoiding catastrophic forgetting that may arise from retraining directly on the target knowledge. Existing methods (Mitchell et al., 2022; De Cao et al., 2021; Zheng et al., 2023; Meng et al., 2022) adopt various strategies: using a hypernetwork to predict parameter updates for incorporating new knowledge (De Cao et al., 2021; Mitchell et al., 2022), identifying and adjusting neurons associated with specific knowledge (Meng et al., 2022, 2023; Jiang et al., 2025), or leveraging in-context learning (ICL) to enforce updated knowledge (Zheng et al., 2023). Beyond correcting factual knowledge, these techniques have also been applied to bias mitigation (Chen et al., 2024a), detoxification (Wang et al., 2024a), personalization (Lu et al., 2025d), unlearning (Li et al., 2025), etc.
More recently, researchers have begun to investigate factual knowledge editing in large vision-language models (LVLMs) (Liu et al., 2023; Li et al., 2023). For example, Cheng et al. (2023) introduced the MMEdit benchmark to explore editing visual factual knowledge, while subsequent works like VLKEB (Huang et al., 2024) and MC-MKE (Zhang et al., 2025) expanded the evaluation scope to provide a more comprehensive understanding of editing in visual modalities. However, no prior work has examined editing auditory attribute knowledge in LALMs, which involves abstract perceptual and auditory concepts rather than concrete facts, distinguishing SAKE from existing research.
@@ 2.2 Large Audio-Language Models (LALMs)
LALMs extend text-based LLMs to auditory modalities such as speech and audio, opening new possibilities for auditory understanding (Gong et al., 2023; Tang et al., 2024; Chu et al., 2023, 2024; Lu et al., 2025b; Ghosh et al., 2025a). These models typically integrate auditory encoders (Radford et al., 2023) with an LLM backbone (Touvron et al., 2023; Yang et al., 2024a) through fine-tuning. While these works advance the integration of auditory knowledge into LLMs, little attention has been given to how auditory-specific knowledge can be edited, which motivates our study.
@@ 3.1 Problem Formulation
We focus on editing auditory attribute knowledge within LALMs, including their perception and understanding of speaker gender, emotion, spoken language, and animal sounds. Here, $y_{e}$ corresponds to new auditory attribute labels (e.g., emotions or languages) that differ from the original attribute labels $y_{o}$ associated with $a_{e}$. For example, given a speech labeled with a happy emotion, we may edit the LALM so that it instead perceives the speech as sad.
@@ 3.2 Evaluation Dimensions and Corresponding Metrics
The edited models should not only generate the correct edit target for the editing data itself but also produce consistent outputs for equivalent neighborhoods of the editing data, such as speech samples sharing the same emotion as $a_{e}$ or paraphrased variants of $x_{e}$. This requirement is quantified by the generality metric $S_{gen}$:
@@
$S_{loc}=\mathbb{E}_{\begin{subarray}{c}(a_{e},x_{e},y_{e})\sim\mathcal{D}_{edit}\\ (a_{\ell},x_{\ell},y_{\ell})\sim\mathcal{L}(a_{e},x_{e},y_{e})\end{subarray}}\Bigl[\mathbb{I}\bigl(f(a_{\ell},x_{\ell};\theta^{\prime})$ (3)
$=f(a_{\ell},x_{\ell};\theta)\bigr)\Bigr].$
@@
Knowledge is not completely disentangled or isolated but rather interconnected. Editing one piece of knowledge may influence other related one. For example, if we edit an LALM’s perception of a frog’s sound to that of a dog, the model’s knowledge of the corresponding physical characteristics of that animal should also be updated. The portability metric $S_{port}$ evaluates how well the editing generalizes the updated knowledge to other related knowledge:
@@
$\mathcal{P}(a_{e},x_{e},y_{e})$ denotes the set of data connected to the edited knowledge. $a_{p}$, $x_{p}$, and $y_{p}$ represent the auditory input, text input, and ground-truth labels of connected instances.
All the alignment and consistency in the definition of these metrics are determined with LLM-as-a-judge (Chiang and Lee, 2023), which is detailed in Sec. 4.3 and Appendix E.
@@ 3.3 Dataset Construction
Auditory Attributes and Audio Sources.
SAKE focuses on various auditory attributes: speaker gender, speaker emotion, spoken language, and animal sound. These attributes are fundamental to speech and audio understanding, widely used in real-world applications, and consistently evaluated in LALM benchmarks (Huang et al., 2025; Sakshi et al., 2025; Yang et al., 2025b). We source audios and attribute labels from SAKURA benchmark (Yang et al., 2025b).
Editing Pairs and Reliability Dataset.
For each attribute, we construct editing pairs $(y_{o},y_{e})$ by uniformly sampling distinct original and target labels, resulting in 300 balanced editing pairs per attribute. To avoid bias, we ensure that all labels for a given attribute appear with approximately equal frequency as both $y_{o}$ and $y_{e}$. Each pair defines an editing instance $(a_{e},x_{e},y_{e})$, where $a_{e}$ is an audio sample labeled with $y_{o}$ and $x_{e}$ is a corresponding attribute recognition question from SAKURA. In total, we construct 1,200 editing instances, forming the editing dataset $\mathcal{D}_{edit}$, which is used both to apply edits and to evaluate reliability.
Generality Dataset Construction.
@@
For each editing instance $(a_{e},x_{e},y_{e})$ in $\mathcal{D}_{edit}$, we construct its equivalent neighborhood $\mathcal{N}(a_{e},x_{e})$ by sampling an alternative audio $a_{e}^{\prime}$ with the same attribute label as $a_{e}$ from the dataset and by paraphrasing the text question $x_{e}$ into $x_{e}^{\prime}$. The paraphrased data are manually verified. Based on these variations, we create testing instances for evaluating generality, considering three cases: (1) Type 1: Textual equivalent neighborhood $(a_{e},x_{e}^{\prime})$; (2) Type 2: Auditory equivalent neighborhood $(a_{e}^{\prime},x_{e})$; (3) Type 3: Equivalent neighborhood involving both auditory and text modalities $(a_{e}^{\prime},x_{e}^{\prime})$. By incorporating these types of testing data, we comprehensively assess how well the editing methods extend the edited knowledge across the equivalent neighborhood.
Locality Dataset Construction.
For each editing instance $(a_{e},x_{e},y_{e})$ in $\mathcal{D}_{edit}$, we build an out-of-scope set $\mathcal{L}(a_{e},x_{e},y_{e})={(a_{\ell},x_{\ell},y_{\ell})}$.
@@
When $a_{\ell}\neq\text{None}$, we consider four types of auditory knowledge locality (Audio Locality): (1) Type 1: Editing one attribute should not affect others. We sample SAKURA question–answer (QA) pairs requiring recognition of attributes different from the edited one; (2) Type 2: Within the same attribute, knowledge regarding labels not involved in the edit (neither $y_{o}$ nor $y_{e}$) should remain unchanged. We sample QA pairs targeting the same attribute with ground truth different from both $y_{o}$ and $y_{e}$, except for gender due to its binary labels; (3) Type 3: When editing from $y_{o}$ to $y_{e}$, the model’s original knowledge of $y_{e}$ should be preserved. We evaluate this using QA pairs with the same attribute and ground truth $y_{e}$; (4) Type 4: Editing should not disrupt general auditory capability. We draw pairs from Dynamic-SUPERB Phase-2 (Huang et al., 2025), excluding tasks involving the edited attributes to ensure irrelevance to the edited knowledge.
When $a_{e}=\text{None}$, we evaluate text locality with textual QA pairs from MMLU (Hendrycks et al., 2021).
Portability Dataset Construction.
@@
For each instance $(a_{e},x_{e},y_{e})$ in $\mathcal{D}_{edit}$, we construct a set of connected knowledge $\mathcal{P}(a_{e},x_{e},y_{e})$ associated with the edited attribute. As SAKURA is designed to evaluate the integration of world knowledge with auditory attributes, it provides attribute-linked knowledge (e.g., physical characteristics of animal labels). Building on this, we design questions that probe $\mathcal{P}(a_{e},x_{e},y_{e})$. To ensure unambiguous evaluation, we exclude questions whose answers are valid for both $y_{o}$ and $y_{e}$ (e.g., avoiding “tail” when editing from dog to cat), ensuring that portability genuinely requires knowledge updating.
Training Dataset Construction and Dataset Summary.
@@
We construct a training dataset for editing methods that require auxiliary training data or access to data beyond test instances. It follows the same procedures as the reliability, generality, and locality datasets, with training and testing sets strictly disjoint to prevent leakage. Statistics are summarized in Table 1, with additional details in Appendix B.2.
Table 1: Summary statistics of SAKE dataset splits.
Split # Instances # Speech/Audio Inputs # QA Pairs
Train 4,000 31,000 35,000
Test 1,200 10,500 11,700